How high is a tree that casts a 26 ft shadow at the same time a 6 ft pole casts a shadow which is 13 ft.
Is it 12 ft

Answer by jim_thompson5910(35256)

(height of tree)/(shadow length of tree) = (height of pole)/(shadow length of pole)

x/26 = 6/13

13x = 26*6

13x = 156

x = 156/13

x = 12

So you are correct.